During year​ 1, Fox Co. started a construction job with a contract price of​ $1,500,000. The job was completed in year 2 and the company uses the percentage of completion method. The following information is available for year 1 and year​ 2:
Year 1 Year 2
Cost incurred to date $350,000 $1,100,000
Estimated cost to complete 650,000 0
Billings to date 300,000 1,100,000
Collections to date 100,000 1,000,000


What amount of gross profit should Fox recognize for this job for year 2?
Group of answer choices
$400,000
$225,000
$325,000
$260,000
Answer #1

Answer:

The  gross profit should Fox recognize for this job for year 2 is $225,000

Working

Year 1 end

Year 2 End

A

Contract Price

$       1,500,000

$       1,500,000

B

Cost Incurred to Date

$           350,000

$       1,100,000

C

Estimated cost yet to be incurred to complete the contract

$       650,000

$                            -  

D = B+C

Total Cost

$       1,000,000

$       1,100,000

E = (B/D) x 100

% of Completion

35.00%

100.00%

F = A x E

Revenue to date

$           525,000

$       1,500,000

G

Revenue of Previous year

$                            -  

$          525,000

H = F - G

Net Revenue this year

$           525,000

$       975,000

I (=B)

Cost to date

$           350,000

$       1,100,000

J

Cost to date of previous year

$                            -  

$          350,000

K = I - J

Net Cost for the year

$           350,000

$       750,000

L = H - K

Gross Profits

$           175,000

$          225,000  Answer